Fermi National Laboratory


Accelerator Update

Wednesday February 11
The day shift began with Operations monitoring store 3226, with the Antiproton Source (Pbar) stacking, with MininBooNE taking beam, and with the Recycler and Main Injector (MI) experts conducting studies.

At 10:17 AM, Recycler experts took three stashes of antiprotons from Pbar that totaled 24mA.

Pbar resumed stacking at 11:19 AM.

From noon to 12:15 PM, Operations noticed a number of I- Source trips. An operator lowered the output current and this helped a little.

MI suffered a major trip of power supplies at 1:18 PM. The cause seemed to be from a NuMI power supply that was turned on and glitched a feeder. Recycler lost its corrector magnet power supplies and lost its stash. The Tevatron (TeV) saw some separator controls problems.

At 3:18 PM, Operations prepared the Recycler for another stash.

Recycler began stashing antiprotons at 6:02 PM.

At 6:12 PM, Pbar resumed stacking.

A Switchyard expert noted a problem with a quadrupole (Q207) at 7:54 PM. They conducted about three and a half hours of studies.

Booster suffered a power supply (EAPS) LCW flow meter alarm at 8:33 PM. Operators found both high and low level alarms. Operations called in an expert who filled the system.

At 11:28 PM, TeV experts began an end-of-store study.

Thursday February 12
The midnight shift began with Pbar stacking, with MiniBooNE taking beam, and with the TeV conducting end-of-store studies.

At 12:28 AM, Operations terminated store 3226.

Operations put the TeV into shot setup at 1:10 AM.

MI had been having problems with their BPMs. At 2:22 AM, an operator rebooted three MI front ends to clear the problem.

At 3 AM, Operation restarted the TeV shot setup. They had found a sequencer problem that left some Septa in the wrong position. Also, the power glitch earlier had changed some settings.

Operations reported that shot setup was still on hold at 3:52 AM due to problems with a MI CAMAC front end. Operations thought the CAMAC problem was tied to the MI BPM problem. Operations called an expert.

At 4:52 AM, Operations reported that the I- Source was still tripping. They tried making some adjustments, but the adjustments didn't help.

Operations established store 3228 at 5:49 AM. Its initial luminosity was 45.7E30.

At 6:05 AM, an Accumulator quadrupole (A:QDF) tripped off, which caused the loss of the stack. Operations contacted an expert.

Operators discovered and put out a small fire on a Debuncher kicker (D:IKIK) high voltage cable at 6:34 AM. It appeared that the cable had shorted out. An expert began repairs.

A Booster expert found a LCW leak on a power supply (EAPS) at 7:34 AM. Water technicians investigated.

An expert completed repairs on D:IKIK at 8:22 AM.

A Linac expert switched to the H- Source and tuned it up at 11:02 AM.

Pbar resumed stacking at 11:46 AM.

CHL notified Operations at 11:51 AM that the cryo system vacuum at TeV sector B3 was going bad. A vacuum pump had tripped off. Vacuum techs will replace this and another vacuum pump at sector C4 after the end of the store.

At 7:38 PM, Operation received a LCW leak alarm for the Booster 95 system.

Operators made Booster ready for an access at 8:30 PM. At 9:08 PM, they discovered a large leak. Operators notified techs; they came in and began repairs at 10:44 PM.

Friday February 13
The midnight shift began with techs working on the Booster LCW leak and with Operations monitoring store 3228.

At 2:11 AM, water techs completed the Booster LCW repairs.

At 4 AM, Operations was trying to clear a LCW flow bit on a Booster power supply (GMPS). The problem seemed to be due to system high pressure. Operations called in a CUB expert.

There were so many devices with water related problems that by 6:30 AM operators had called in may experts.

Operations established beam to MiniBooNE at 7:41 AM.

Pbar resumed stacking at 7:46 AM.

Today's Plans
The plans for today and this weekend are to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/13/2004   email Fermilab